Show each step in solving and explain the meaning of the solutions to the following problems:
4x-3<9x+27

Answer by kendra4227(8)

4x-3<9x+27
4x-3+3<9x+27-27
-5x=27+3
-5x=30
-5x/-5=30/-5
-x=-6
-x<-6
Because -x= -1 so -1<-6